# Relationship between household attributes and contact patterns in urban and rural South Africa

### v3

@@ 2 Results
Figure 2: Contact matrices. Each frame represents one of the aggregations described in Section 2, plus the one obtained by grouping measurement waves. The matrices’ entries denote the daily average contact duration per household by gender-age group, expressed in minutes and normalized by the population sizes, and are obtained following the procedure detailed in Section 4.4. The number and color-code report the average value of the bootstrap, while the numbers in the bracket are the $10\%$ confidence interval. The population sizes for each aggregation level are reported in Table 1. The matrices are indexed by gender (“M for male, and “F for female) and age (“Ch” for children, “Ado” for adolescents, and “Adu” for adults).
@@
Figure 3: Household interaction time per individual. Top row: daily household interaction times normalized by the population, disaggregated by wave, site, household type, and gender of the household head. The $y$-axis indicates the sum of all entries on the upper diagonal of the normalized household contact matrices shown in Figure 2. The distributions are obtained from a bootstrap sample, as described in Section 4.4. Bottom row: ridge regression. Box plot of the ridge regression coefficients obtained from several bootstrap randomizations. The whiskers show the 5955-95 confidence interval. The face color is gray if the confidence interval crosses zero, and it is orange otherwise.
@@ 2.1 Comparison of the interaction rates
We compare the interaction rates per individual for different household attributes. We consider the interaction rates per household, by age and gender, and with children. The analysis is performed on a bootstrapped sample of the empirical contact matrices, which is detailed in Section 4.4. Size effects are quantified in terms of the Cohen’s $d$ score, and larger scores imply higher separations between the distributions. To handle the dependencies between the household attributes (see Figure 1), we run a ridge regression to predict the interaction rates from the household attributes. This analysis allows us to identify statistically significant effects in predicting the interaction rates. The details of the ridge regression implementation are provided in Section 4.5. The statistical analysis is performed in Python, using the numpy, pandas, and scipy packages.
@@ 2.2 Epidemiological impact of age and gender stratification according to the household classification
Multi-class compartmental models, such as the SIR (Susceptible-Infected-Recovered) model, are widely used in epidemiology [54]. In this model, susceptible individuals ($S$) become infected ($I$) when entering into contact with already infected individuals. After some time, infected individuals recover ($R$). By including contact matrices in this model, one can account for across-group heterogeneities in the social mixing behavior. In this model, two parameters – $\beta$ and $\mu$ – quantify the probability per unit time of infection and recovery, respectively. These parameters are disease-dependent and determine the basic reproductive number $R_{0}$, which is the average number of secondary infections caused by a single infected person in a fully susceptible population. Its value crucially determines the epidemic outcome since if $R_{0}<1$, the epidemic likely dies quickly, while if $R_{0}>1$, it is expected to spread rapidly through the population. For the SIR model, the $R_{0}$ has a known expression depending on the disease parameters $\beta,\mu$ and on the contact matrix $R$ [55]:
@@
Figure 6: Largest eigenvalue of the contact matrix for different aggregation and indexing strategies. Each panel shows an aggregation based on a specific attribute indicated by the title. We consider three indexing strategies (shown on the x-axis) for each matrix: age, gender, and age-gender. The $y$-axis shows $\rho(RN^{-1})$, the largest eigenvalue of contact matrices the normalized contact matrix, which is proportional to $R_{0}$. The box plots are obtained for several bootstrap randomizations, as detailed in Section 4.4. The whiskers indicate the 5955-95 confidence interval.
@@
The matrices shown in Figure 2 utilize a two-level stratification that combines age and gender. To evaluate the significance of this stratification, Figure 6 compares the spectral radius of three contact matrices for each of the household classifications introduced in Section 2: one using only gender for indexing groups, another using only age, and a third combining both age and gender. The results are obtained for several bootstrap randomizations, as detailed in Section 4.4. As expected, age accounts for a higher level of heterogeneity than gender, but less than the combination of both attributes. Table 1 reports the size effects between the distributions of $R_{0}$, comparing “Gender”-only and “Age”-only contact matrices with “Age-Gender” matrices, testifying the importance of including both variables to explain the contact patterns in extended households where the complex patterns of interactions described in the previous sections lead to the largest increase in $R_{0}$ when using both variables at once.
@@ 4.2 Data description
The PHIRST study [50, 26, 29] was a perspective household cohort study, conducted in South Africa during 201620182016-2018 to investigate the burden and transmission dynamics of influenza and respiratory syncytial virus (RSV). Randomly selected households were approached for enrolment, and households with more than $2$ members and where at least $80\%$ of household members consented, were included in the study. A new cohort was enrolled each year, with enrolment occurring in the November through to December the prior year, and intensive follow up occurring from February - October with twice weekly home visits and nasopharyngeal swab collection for testing of influenza and (RSV), irrespective of symptoms. Recruitment occurred at the end of the preceding year to ensure all households were enrolled before the start of the RSV season which starts in February/March in South Africa. A household contact survey was nested in PHIRST in 2018, in which all households included in the 2018 cohort were approached to partake. During this data collections, contact patterns were measured with the proximity sensors developed by the SocioPatterns collaboration (sociopatterns.org, [51]). These sensors are non-obtrusive devices worn on the chest of the study participants during the data collection. They emit low-power radio signals, exchanging approximately $40$ information packets per second when facing one another. They record a proximity interaction if two sensors exchange at least one packet in a $20$-second period. In this case, each sensor records the unique identifier of another interacting sensor, the interaction time, and the power attenuation from the sender to the receiver. Proximity interactions within approximately $1.5$ meters are identified by filtering the measurements based on their signal attenuation. These sensors provide high-resolution measurements of human interactions and have been deployed in several contexts, including schools [64, 65], hospitals [66], and conferences [67], among others. The proximity sensors were packaged in PVC pouches and worn as a badge pinned to either the inside or outside of the clothing with a safety pin, or on a lanyard around the neck, depending on the preference of the participant. Field workers visited the household at least two times during the deployment and reminded participants to wear the sensors. We followed the data cleaning procedure described in [29], which we here briefly summarize.
@@ 4.3 Household classification
  • Gender of the household head. The household head does not have a unique definition, but it generally refers to the person who holds the economic responsibility (for example being the primary breadwinner) or who has the most decision-making power in the household [68]. In our study, the household head is the person identified by the household members as the primary decision-maker. Especially in South Africa, this role is commonly held by the oldest individuals who typically also have the highest income [69]. Regardless of the actual dynamics within the household, men are commonly more likely to be designated as household heads [68], while women-headed households are more susceptible to economic hardships [69]. Our dataset comprises $43$ female-headed households ($24$ in Agincourt, $19$ in Klerksdorp) and $17$ male-headed households ($4$ in Agincourt and $13$ in Klerksdorp). This agrees with the general trend in South Africa that sees female-headed households more prevalent in rural areas [70].

### v4

@@ 2 Results
Figure 2: Contact matrices. Each frame represents one of the aggregations described in Section 2, plus the one obtained by grouping measurement waves. The matrices’ entries denote the daily average contact duration per household by gender-age group, expressed in minutes and normalized by the population sizes, and are obtained following the procedure detailed in Section $4.$ The number and color-code report the average value of the bootstrap, while the numbers in the bracket are the $10\%$ confidence interval. The population sizes for each aggregation level are reported in Table 1. The matrices are indexed by gender (“M for male, and “F for female) and age (“Ch” for children, “Ado” for adolescents, and “Adu” for adults).
@@
Figure 3: Household interaction time per individual. Top row: daily household interaction times normalized by the population, disaggregated by wave, site, household type, and gender of the household head. The $y$-axis indicates the sum of all entries on the upper diagonal of the normalized household contact matrices shown in Figure 2. The distributions are obtained from a bootstrap sample, as described in Section $4.$ Bottom row: ridge regression. Box plot of the ridge regression coefficients obtained from several bootstrap randomizations. The whiskers show the 5955-95 confidence interval. The face color is gray if the confidence interval crosses zero, and it is orange otherwise.
@@ 2.1 Comparison of the interaction rates
We compare the interaction rates per individual for different household attributes. We consider the interaction rates per household, by age and gender, and with children. The analysis is performed on a bootstrapped sample of the empirical contact matrices, which is detailed in Section $4.$ Size effects are quantified in terms of the Cohen’s $d$ score, and larger scores imply higher separations between the distributions. To handle the dependencies between the household attributes (see Figure 1), we run a ridge regression to predict the interaction rates from the household attributes. This analysis allows us to identify statistically significant effects in predicting the interaction rates. The details of the ridge regression implementation are provided in Section 4.5. The statistical analysis is performed in Python, using the numpy, pandas, and scipy packages.
@@ 2.2 Epidemiological impact of age and gender stratification according to the household classification
Multi-class compartmental models, such as the SIR (Susceptible-Infected-Recovered) model, are widely used in epidemiology [54]. In this model, susceptible individuals ($S$) become infected (II) when entering into contact with already infected individuals. After some time, infected individuals recover ($R$). By including contact matrices in this model, one can account for across-group heterogeneities in the social mixing behavior. In this model, two parameters – $\beta$ and $\mu$ – quantify the probability per unit time of infection and recovery, respectively. These parameters are disease-dependent and determine the basic reproductive number $R_{0}$, which is the average number of secondary infections caused by a single infected person in a fully susceptible population. Its value crucially determines the epidemic outcome since if $R_{0}<1$, the epidemic likely dies quickly, while if $R_{0}>1$, it is expected to spread rapidly through the population. For the SIR model, the $R_{0}$ has a known expression depending on the disease parameters $\beta,\mu$ and on the contact matrix $R$ [55]:
@@
Figure 6: Largest eigenvalue of the contact matrix for different aggregation and indexing strategies. Each panel shows an aggregation based on a specific attribute indicated by the title. We consider three indexing strategies (shown on the x-axis) for each matrix: age, gender, and age-gender. The $y$-axis shows $\rho(RN^{-1})$, the largest eigenvalue of contact matrices the normalized contact matrix, which is proportional to $R_{0}$. The box plots are obtained for several bootstrap randomizations, as detailed in Section $4.$ The whiskers indicate the 5955-95 confidence interval.
@@
The matrices shown in Figure 2 utilize a two-level stratification that combines age and gender. To evaluate the significance of this stratification, Figure 6 compares the spectral radius of three contact matrices for each of the household classifications introduced in Section 2: one using only gender for indexing groups, another using only age, and a third combining both age and gender. The results are obtained for several bootstrap randomizations, as detailed in Section $4.$ As expected, age accounts for a higher level of heterogeneity than gender, but less than the combination of both attributes. Table 1 reports the size effects between the distributions of $R_{0}$, comparing “Gender”-only and “Age”-only contact matrices with “Age-Gender” matrices, testifying the importance of including both variables to explain the contact patterns in extended households where the complex patterns of interactions described in the previous sections lead to the largest increase in $R_{0}$ when using both variables at once.
@@ 4.2 Data description
The PHIRST study [50, 26, 29] was a perspective household cohort study, conducted in South Africa during 201620182016-2018 to investigate the burden and transmission dynamics of influenza and respiratory syncytial virus (RSV). Randomly selected households were approached for enrolment, and households with more than 22 members and where at least $80\%$ of household members consented, were included in the study. A new cohort was enrolled each year, with enrolment occurring in the November through to December the prior year, and intensive follow up occurring from February - October with twice weekly home visits and nasopharyngeal swab collection for testing of influenza and (RSV), irrespective of symptoms. Recruitment occurred at the end of the preceding year to ensure all households were enrolled before the start of the RSV season which starts in February/March in South Africa. A household contact survey was nested in PHIRST in 2018, in which all households included in the 2018 cohort were approached to partake. During this data collections, contact patterns were measured with the proximity sensors developed by the SocioPatterns collaboration (sociopatterns.org, [51]). These sensors are non-obtrusive devices worn on the chest of the study participants during the data collection. They emit low-power radio signals, exchanging approximately $40$ information packets per second when facing one another. They record a proximity interaction if two sensors exchange at least one packet in a 2020-second period. In this case, each sensor records the unique identifier of another interacting sensor, the interaction time, and the power attenuation from the sender to the receiver. Proximity interactions within approximately $1.5$ meters are identified by filtering the measurements based on their signal attenuation. These sensors provide high-resolution measurements of human interactions and have been deployed in several contexts, including schools [64, 65], hospitals [66], and conferences [67], among others. The proximity sensors were packaged in PVC pouches and worn as a badge pinned to either the inside or outside of the clothing with a safety pin, or on a lanyard around the neck, depending on the preference of the participant. Field workers visited the household at least two times during the deployment and reminded participants to wear the sensors. We followed the data cleaning procedure described in [29], which we here briefly summarize.
@@ 4.3 Household classification
  • Gender of the household head. The household head does not have a unique definition, but it generally refers to the person who holds the economic responsibility (for example being the primary breadwinner) or who has the most decision-making power in the household [68]. In our study, the household head is the person identified by the household members as the primary decision-maker. Especially in South Africa, this role is commonly held by the oldest individuals who typically also have the highest income [69]. Regardless of the actual dynamics within the household, men are commonly more likely to be designated as household heads [68], while women-headed households are more susceptible to economic hardships [69]. Our dataset comprises $43$ female-headed households ($24$ in Agincourt, 1919 in Klerksdorp) and $17$ male-headed households (44 in Agincourt and $13$ in Klerksdorp). This agrees with the general trend in South Africa that sees female-headed households more prevalent in rural areas [70].